Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis

Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 78,000 units of product were as follows:


Standard CostsActual Costs
Direct materials226,200 lbs. at $5.30223,900 lbs. at $5.10
Direct labor19,500 hrs. at $16.6019,950 hrs. at $17.00
Factory overheadRates per direct labor hr.,

based on 100% of normal

capacity of 20,350 direct

labor hrs.:


Variable cost, $4.30$83,010 variable cost


Fixed cost, $6.80$138,380 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Material Price Variance$
Direct Materials Quantity Variance$
Total Direct Materials Cost Variance$

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct Labor Rate Variance$
Direct Labor Time Variance$
Total Direct Labor Cost Variance$

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance$
Fixed factory overhead volume variance$
Total factory overhead cost variance$
Answer #1
a)
D.mat price variance= AQ*SP-AQ*AP
D.mat price variance= 223900*5.30-223900*5.10
D.mat price variance= $                                                                                                  (44,780) Favourable
D.mat quantity variance= SQ*SP-AQ*SP
D.mat quantity variance= 226,200*5.30-223900*5.30
D.mat quantity variance= $                                                                                            (12,190.00) Favourable
Total Direct Materials Cost Variance= SQ*SP-AQ*AP
Total Direct Materials Cost Variance= 226,200*5.30-223900*5.10
Total Direct Materials Cost Variance= $                                                                                            (56,970.00) Favourable
b)
D.LABOR RATE variance= AH*SR-AH*AR
D.LABOR RATE variance= 19,950*16.60-19,950*17.00
D.LABOR RATE variance= $                                                                                                      7,980 UnFavourable
D.LABOR TIME variance= SH*SR-AH*SR
D.LABOR TIME variance= 19,500*16.60-19,950*16.60
D.LABOR TIME variance= $                                                                                                      7,470 UnFavourable
Total Direct LABOR Cost Variance= SH*SR-AH*AR
Total Direct LABOR Cost Variance= 19,500*16.60-19,950*17.00
Total Direct LABOR Cost Variance= $                                                                                                    15,450 UnFavourable
c)
Variable factory OH controllable variance= Actual VOH-Budgeted VOH
Variable factory OH controllable variance= 83,010-19,500*4.30
Variable factory OH controllable variance= $                                                                                                        (840) Favourable
Fixed OH volume variance= Absorbed Fixed OH-Budgeted Fixed OH
Fixed OH volume variance= Actual units*Absorbtion rate-Budgeted units*absorbtion rate
Absorbtion rate= Fixed cost per hour*0.25
Absorbtion rate= 6.80*0.25= $ 1.70
Fixed OH volume variance= 78000*1.70-78000*1.70
Fixed OH volume variance= 0
No Fixed OH volume variance exists because actual output = budgeted output
Total factory overhead cost variance= Total Variable OH variance + Total Fixed OH variance
Total Variable OH variance= Actual Variable OH - budgeted Variable OH
Total Variable OH variance= 83010-19,500*4.30
Total Variable OH variance= $                                                                                                        (840) Favourable
Total Fixed OH variance= Actual Fixed OH-Absorbed Fixed OH
Total Fixed OH variance= 138380-78000*1.70
Total Fixed OH variance= $                                                                                                      5,780 UnFavourable
Total factory overhead cost variance= $                                                                                                      4,940 UnFavourable